Suppose that $.90 U.S. = 1 Canadian dollar, and .60 British pounds = $1 U.S. What is the cross rate of the Canadian dollar for the British pound? How many Canadian dollars for a pound?
Answer #1

Canadian Dollar/US = 0.9

US/British Pound = 0.6

Canadian Dollar/British Pound = Canadian Dollar/US * US/British Pound = 0.9*0.6 = 0.54

1 Canadian Dollar = 0.54 British Pound

1 British Pound = 1/0.54 Canadian Dollar = 1.852 Canadian Dollar